OpenStudy (anonymous):

how do i solve for |(1+3x)/5|>7/8

OpenStudy (anonymous):

\[|\frac{1+3x}{5}|>\frac{7}{8}\]?

OpenStudy (dumbsearch2):

@satellite73, I have a quick question :3 You may dismiss it as stupid, but... are you hired by OpenStudy ?

OpenStudy (anonymous):

multiply both sides by 5 as a first step get \[|1+3x|>\frac{35}{8}\]

OpenStudy (anonymous):

@dumbsearch2 no, i am free lance

OpenStudy (anonymous):

but i do charge $10 for answers $20 for correct ones

OpenStudy (dumbsearch2):

Just, I just don't understand how anyone would devote so much of their time and intelligence to answer 23375 questions!!

OpenStudy (dumbsearch2):

haha your joking tho, right ? :P

OpenStudy (dumbsearch2):

"$20 for correct ones" LOL.

OpenStudy (anonymous):

\[1+3x>\frac{35}{8}\] or \[1+3x<-\frac{35}{8}\]

OpenStudy (anonymous):

yes satellite>7/8

OpenStudy (dumbsearch2):

IDK, You just don't seem like a regular user. You don't comment on the feedback section, you've only become a mod recently (really weird! i mean, with 100 SS, and all those amazing 15716 medals, I'd expect you'd become one sooner), you don't linger much in chat, and answering so many questions just seem surreal, that you would so much of your time and intelligence to answer 23375 questions!!

OpenStudy (dumbsearch2):

@satellite73 Sorry, though I really thought you were on a paycheck! :|